What is the first approximation of the function f(x)=(1−x)^100 near x=0?
1
Expert's answer
2020-05-08T20:18:39-0400

It is necessary to find the first approximation of the function "f(x) = (1-x)^{100}" near x=0?


The first approximation of the function is a line with the equation "y- y_0=k(x-x_0)," where "x_0=0" by the condition. "y_0=f(x_0 )=(1-0)^{100}=1."

To find k we must find the first derivative of f(x) and substitute "x_0=0". So:

"\frac{df(x)}{dx}=-100(1-x)^{99}"

Then:

"k=\frac{df(x_0)}{dx}=-100(1-0)^{99}=-100"

Write an equation of the straight line:

"y-1=-100x"

Answer: the first approximation of the function "f(x) = (1-x)^{100}" is the line "y=-100x+1".